### v2

# OmniGenome: Aligning RNA Sequences with Secondary Structures in Genomic Foundation Models

Abstract: The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated. Existing FMs have struggled to establish sequence-structure alignment, hindering the free flow of genomic information between RNA sequences and structures. In this study, we introduce OmniGenome, an RNA FM trained to align RNA sequences with respect to secondary structures based on structure-contextualised modelling. The alignment enables free and bidirectional mappings between sequences and structures by utilising the flexible RNA modelling paradigm that supports versatile input and output modalities, i.e., sequence and/or structure as input/output. We implement RNA design and zero-shot secondary structure prediction as case studies to evaluate the Seq2Str and Str2Seq mapping capacity of OmniGenome. Results on the EternaV2 benchmark show that OmniGenome solved of puzzles, whereas existing FMs only solved up to of the puzzles due to the oversight of sequence-structure alignment. We leverage four comprehensive in-silico genome modelling benchmarks to evaluate performance across a diverse set of genome downstream tasks, where the results show that OmniGenome achieves state-of-the-art performance on RNA and DNA benchmarks, even without any training on DNA genomes.
Keywords: RNA Foundation Model, RNA Design, Secondary Structure Prediction
1 Introduction
RNA is a critical type of molecule that encodes a vast array of biological regulatory elements that orchestrate crucial aspects of plant growth, development, and adaptation to environmental stresses. To decipher the genomic code in RNA and manipulate RNA engineering and design, current research mainly uses bioinformatics in solving RNA genome-oriented challenges. Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ [1] hidden in genome encoding processes. Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction [2, 3, 4, 5], degradation rate prediction [6, 7], and mRNA vaccine design [8, 9]. In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology [10]. However, the role of the structure as a second ‘genomic language’ to interact with sequences and solve various RNA downstream tasks has been largely ignored.
Sequence-Structure Alignment in FMs
We define alignment between sequences and secondary structures111In this paper, all structures are referred to as the secondary structures. as the bidirectional information flows. Current FMs have been struggling to establish an alignment between RNA nucleotide sequences and their folded structures, thus impeding bidirectional genomic information flows. There has been a deep scientific challenge to align RNA sequences with structures because it is not deterministic to predict sequences from structures and vice versa. In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors [11]. Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Fig. 1. The oversight of such alignment in existing FMs causes outstanding issues in understanding and leveraging RNA structures, such as mRNA design. For example, recent state-of-the-art RNA FMs, RNA-FM [12] and RNA-MSM [13], only solved out of puzzles in in-silico RNA design [14]. This is because they fail to decipher corresponding sequences based on structures to guide RNA design.
To address the above two problems, we propose sequence-structure alignment in RNA FMs, which leverages the large-scale annotations of sequences and structures to build reliable structure to sequence (Str2Seq) and sequence to structure (Seq2Str) mappings, leading to an aligned FM dubbed OmniGenome. The sequence-structure alignment enables genomic information to freely flow between sequences and structures by introducing a flexible RNA modelling paradigm that supports versatile inputs and outputs modalities, i.e., sequence and/or structure as input/output. The sequence-structure alignment enables genomics information to freely flow between sequences and structures by introducing a flexible RNA modelling paradigm that supports versatile inputs and outputs modalities, i.e., sequence and/or structure as input/output. Furthermore, the sequence-structure alignment is designed to be architecture-agnostic and genome-agnostic. That is to say, it can be easily transferred to large-scale models with new architecture and different genome types like DNA.
Str2Seq Mapping
RNA structure serves as a vital input in most of the RNA genome analysis tasks. To induce the ability of Str2Seq mapping in genomic FMs, we formulate a structure-contextualised RNA sequence reconstruction task, which stems from the representation of RNA secondary structures in texts composed of dots and brackets. As the diagram is shown in Fig. 2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence. Then, we pre-train OmniGenome to reconstruct the masked nucleotide bases given the structure contexts. This simple but effective formulation of Str2Seq mapping realises structure input awareness in genomics FM pre-training and provides substantial compatibility for structure-contextualised tasks, which has been verified in our RNA design benchmark.
Seq2Str Mapping
On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) [15, 16], is another critical aspect of achieving the alignment. We generalise end-to-end structure pre-training [17] to OmniGenome pre-training. This large-scale structure pre-training on diversified genomes supervises OmniGenome to perform Seq2Str mapping. The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions [18] and barriers the structure prediction robustness on small datasets. To conduct Seq2Str mapping, tremendous secondary structure annotations are required to avoid data bias. A feasible solution to RNA structure pre-training is leveraging the plausible structures calculated based on the minimum free energy. In this paper, we leverage the popular ViennaRNA [19] to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
Evaluation and Results
To validate the effectiveness of OmniGenome, we designed four large-scale genome benchmarks with diverse genomics tasks. The first one is the RNA genomics benchmark (RGB) compiled in the study, which contains diverse challenging genomics understanding tasks that benefit from the sequence-structure alignment, such as degradation rate prediction. The second benchmark is the plant genomics benchmark (PGB) [20] which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks. In particular, we want to use this benchmark to evaluate the generalisability of OmniGenome among diversified species and genomes. The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT [20] which contains billion parameters. The last two benchmarks, available in the appendix, are the genomics benchmark (GB) [21] and genomics understanding evaluation (GUE) [22], which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
In addition, we also conduct zero-shot Seq2Str and Str2Seq prediction experiments to verify the performance of sequence-structure alignment. As revealed in the experiments in Sections 3.3 and 3.3, OmniGenome achieves up to a macro-F1 score in zero-shot Seq2Str prediction, i.e., secondary structure prediction, outperforming fine-tuned FMs and bioinformatics methods like ViennaRNA. In terms of Str2Seq prediction performance, we evaluate the performance of OmniGenome in the in-silico RNA design task. We solved of complex puzzles of the EternaV2 benchmark [14], while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to . Besides, OmniGenome only takes less than one hour to solve most of the puzzles, while most RNA design methods need to take up to hours to solve even a single puzzle.
Open-source Toolkit and Tutorials
Open science is always the golden standard to promote this rising area of FM for genome modelling, which unfortunately lacks relevant high-quality resources such as code integrity, data availability, and pre-training pipeline. To address this gap, following the FAIR principles [23], we developed an open-source package222https://tinyurl.com/OmniGenome that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few. It provides ready-to-use genomics benchmarks and uses the API with only a few lines of code to streamline benchmarking purposes. We believe this will be a valuable resource to make this emerging AI for the RNA community to thrive.
2 Methodology
This section delineates the implementation details of OmniGenome including its entire pre-training workflow and downstream benchmarks.
2.1 RNA Tokenization for Alignment
We aim to implement a fine-grained alignment between RNA sequences and structures, where each base in the sequences reflects a structural label in ‘(’, ‘)’, ‘.’. Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method [1, 24] in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels. We illustrate our tokenization based on an example shown in Fig. 3.
Our adapted SNT features bidirectional mappings between single nucleotide (SN) bases and structural labels required by sequence-structure alignment. Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution [1, 24]. For example, the k-mers [25, 26] and BPE [27, 22] tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction. Like other encoder-only models, e..g, BERT [27], we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
2.2 Pre-training Objectives
As discussed in Section 1, a key desideratum for SN-level genome modelling is to build the alignment between RNA sequences with corresponding secondary structures. Bearing this in mind, we formulate two pre-training objectives, i.e., and , for Str2Seq and Seq2Str predictions, respectively. Besides, we aggregate these two objectives with the masked RNA language modelling objective MRLM to pre-train OmniGenome as follows:
where is the regularisation weight and represents the parameters of OmniGenome. The following paragraphs explain the design principles of each objective function used in equation (1).
is designed to enable OmniGenome to predict bases given structure-contextualised sequences with partially masked bases. This objective aims at Str2Seq tasks and teaches OmniGenome to interpret structure information and infer the masked sequences. To achieve this objective, we mask of the bases and structure tokens, encouraging the model to infer masked bases (i.e., ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’) and structure tokens (i.e., ‘(’, ‘)’, ‘.’). Specifically, is defined as the classic cross-entropy loss widely used in the masked language modelling:
(2) where is the number of masked nucleotide and structure tokens, and indicates the probability of predicting the masked nucleotide based on its context.
In terms of structure-out modelling, we implement to enable OmniGenome for Seq2Str predictions. Instead of directly feeding the secondary structure into OmniGenome as inputs, this objective employs the RNA secondary structures as labels for supervised training. This objective is implemented as a token-level classification, where the loss is defined in the following cross-entropy loss:
(3) where denotes the label of secondary structure at the -th position, and is the probability predicted by a linear classifier deployed on OmniGenome. is the length of an RNA sequence and denotes the number of the possible labels of structure, i.e., ‘(’, ‘)’, ‘.’.
The last objective, , is adapted to the conventional masked language modelling loss in NLP. It aims to improve the model’s understanding of genomic language in RNA sequences by predicting the masked or replaced of nucleotide bases. The definition of is similar to that of which only considers the prediction of masked bases rather than randomly replaced bases. The loss function of MRLM is well-known so we omit its formula here.
We cannot trust structure predictions (in ) while the structures are leaked in inputs (in ), i.e., the sequence inputs and outputs of these two objectives are exclusive. In practice, we only consider objectives either or for each input sequence. In the pre-training, of RNA sequences are used for the first two objectives, while the remaining are used for the latter two objectives. This proportion setting is concluded from our empirical experience to balance the capability of Str2Seq and Seq2Str predictions.
2.3 Model Architecture
OmniGenome adopts the Transformer encoder architecture with bidirectional multi-head attention. We do not adopt recent architectures like Mamba [28, 29] and Hyena [1] because our experiments in Table 5 and Table 6 show that these architectures are not competent at RNA genome understanding. This low performance is probably because RNA sequences are much shorter than DNA sequences in the wild.
We designed two variants, dubbed OmniGenome and OmniGenome with and million parameters respectively. Some key model specifications are summarised in Table 1.
| OmniGenome | 52M | 186M |
| # of Layers | ||
| Embedding dimension | ||
| Intermediate dimension | ||
| # of heads | ||
| # of parameters | M | M |
| Modelling length | ||
To improve the reproducibility of OmniGenome, we list the pre-training settings and hyperparameters as follows.
The learning rate is set to and the weight decay is set to .
We use AdamW as the optimiser with hyperparameters and .
We use a linear decay strategy with a warm-up period of steps in the learning rate scheduler.
The batch size is set to .
No dropout is applied during pre-training, and we use the rotary position embeddings [30] to further enhance the model’s scalability to long RNA sequences.
We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix 6. The pre-training was finished in approximately 1 and 3 weeks for OmniGenome and OmniGenome, respectively.
2.4 Pre-training Database: OneKP
Recent studies [24, 22] have shown that data diversity can enhance FM performance without significantly increasing model capacity. For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ [31], which compiles large-scale RNA raw sequence database from plant species. The raw sequences are not available for pre-training before processing and filtering.
We adopt the following raw sequence data curation protocol to fit pre-training.
To enhance training efficiency and reduce bias, we removed all duplicate sequences.
To tackle incomplete transcriptome data and other noises, we discard sequences shorter than bases.
To facilitate the sequence-structure alignment training, we adopt ViennaRNA444https://github.com/ViennaRNA/ViennaRNA to obtain the secondary structures for the sequences.
2.5 Benchmark Suites
OmniGenome is designed as a general-purpose RNA FM that can be fine-tuned for a diverse set of downstream genomics predictive tasks. In this paper, we constructed a large-scale benchmark suite for RNA FMs. According to the category of genomes, we split the benchmark into two parts.
RNA Genomic Benchmark (RGB)
RGB is a collection of genome understanding tasks, as shown in Table 7. RGB contains SN-level tasks that are curated in this work or collected from published articles. The purpose of RGB is to benchmark FMs in challenging SN-level modelling tasks like the predictions of mRNA degradation rates and secondary structures. The sequence length in RGB ranges from to , which is enough for most RNA understanding tasks. These multi-species and SN-level tasks in RGB serve as the first comprehensive RNA benchmark to assess the modelling capabilities of FMs. For detailed information on each dataset, such as their sources and sizes, please refer to Appendix 9.1.
Plant Genomic Benchmark (PGB)
PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark [20] ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology. PGB involves types of DNA downstream subtasks, including a range of critical tasks such as promoter strength prediction and gene expression regression. There are datasets in total with millions of DNA sequences to be evaluated in PGB, and the sequence lengths are up to , which is very long for most of the genomic FMs. Since the original evaluation protocol is not publicly available, we have re-implemented the auto-benchmark for all the subtasks from PGB in our package. By integrating diverse downstream tasks, PGB aims to facilitate the development of plant genomics and robust assessment. Due to computational limitations, we randomly sample a maximum of k examples in all downstream datasets in PGB to evaluate the FM’s performance.
2.6 Str2Seq Modelling Case: RNA Design
One of the most challenging practices addressed by OmniGenome is RNA design, which has not been settled in existing FMs because of the oversight of sequence-structure alignment. In this section, we provide a case of RNA design that requires the exploitation of sequence-structure alignment. To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/OmniGenome/tutorials/zeroshot_omnigenome_RNA_design.py to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome. The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix 10 and Fig. 5, respectively. In Section 3.3, the experimental results on the Eterna V2 dataset indicate an impressive performance in RNA design compared to existing methods.
3 Experiments
To evaluate the performance of OmniGenome across genome modelling, we implement experiments on diverse downstream tasks. We first evaluate the sequence-structure alignment capability of OmniGenome. Subsequently, we evaluate the overall performance of OmniGenome on two comprehensive genomic modelling benchmarks, i.e., RGB and the PGB, respectively. Finally, we include the GB and GUE in the appendix to evaluate the performance on non-plant genomes.
3.1 RNA Sequence Filtering
The pertaining involves RNA sequences and structures prediction, we take the data and annotation leakage problem seriously.
To avoid structure annotation leakage of downstream benchmarks, the secondary structure predictors for all FMs were randomly initialised for fair comparisons, which means the pre-trained structure predictor of OmniGenome was not used in benchmarks, except for zero-shot SSP experiments. Please find the source codes for details.
To reduce sequence leakage caused by evolutionary conservative sequences across multiple species, we use the ch-hit-est tool to calculate the sequence similarity between sequences from the OneKP database and downstream tasks. We adopt the similarity threshold of for ch-hit-est to eliminate sequences whose homogeneous sequences appeared in the OneKP database. Subsequently, we exploit the blastn tool to query potentially leaked sequences in downstream benchmark datasets and further alleviate the data leakage problem. The e-value has been set to for rigorous sequence filtering.
| Model | Tokenization | # of Params | Pre-training Data Scale | Pre-training Data Source | Species | Sequence Type |
| DNABERT-2 | BPE | M | B Tokens | The 1000 Genomes Project | Human + Species | DNA |
| NT-V2-M | k-mers | M | B Tokens | The 1000 Genomes Project, etc. | Human + Species | DNA |
| HyenaDNA-Large | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
| Caduceus | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
| Agro-NT-B | k-mers | M | B Tokens | Ensembl Plants Database | Edible Plants | DNA |
| SpliceBERT | SNT | M | M Sequences | UCSC Genome Browser | Multi-Vertebrates | precursor-mRNA |
| RNA-BERT | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
| RNA-MSM | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
| RNA-FM | SNT | M | M Sequences | RNAcentral Database | Multi-Species | ncRNA |
| 3UTRBERT | k-mers | M | Sequences | The GENCODE Project | Human | mRNA ’UTR |
| OmniGenome | SNT | M | B Tokens | The OneKP Initiative | Plant Species | mRNA, CDS, UTR |
| OmniGenome | M |
3.2 Comparison Baselines
Apart from OmniGenome, we implement a plus variant, i.e., OmniGenome. In the context of OmniGenome, we assume the structure annotation from ViennaRNA is always available for enhancing the model based on structure-contextualised modelling. In SSP tasks, we can also use the ViennaRNA’s structure annotations as contexts to improve downstream SSP performance. Please refer to Appendix 8 for brief introductions of these FMs.
We can compare OmniGenome with the following RNA and DNA FMs shown in Table 2 as baselines to help evaluate the performance of OmniGenome. We are aware that some FMs are also developed for RNA, such as Uni-RNA [34], 5UTR-LM [35], etc. However, we cannot compare OmniGenome with them because their source codes are very hard to work with in our efforts or not publicly available.
3.3 Sequence-Structure Alignment Evaluation
In this section, we verify the sequence-structure alignment capability based on two experiments, i.e., Str2Seq prediction and zero-shot Seq2Str prediction via SSP and RNA design tasks, respectively. Overall, the results in Table 3 and Table 4 provide reliable evaluations of the FMs’ capabilities in sequence-structure alignment. This underscores OmniGenome’s efficacy in enabling genomic information to freely flow between structures and sequences.
RNA Design (Str2Seq) Evaluation
we demonstrate the Str2Seq prediction capability of OmniGenome based on RNA design. We employed the Eterna777https://github.com/eternagame/eterna100-benchmarking/blob/master/eterna100_vienna2.txt [14] V2 benchmark, which consists of specified secondary structures. This task aims to design RNA sequences based on reference structures. We develop a genetic algorithm (GA) which exploits masked nucleotide modelling (a.k.a., masked language modelling) to find plausible RNA sequences that solve RNA design puzzles. The implementation details can be found in Fig. 5 in Appendix 10. In the GA, the population size is set at , with iterations, and the mutation rate for each base is . The evaluation metric is accuracy following existing works which indicates the number of puzzles solved by FMs. The experimental results are available in Table 3.
| Model | Token. | EternaV2 (Acc) |
| RNAInverse | — | |
| 3UTRBERT | k-mers | |
| DNABERT2 | BPE | |
| SpliceBERT | SNT | |
| RNA-MSM | SNT | |
| RNA-FM | SNT | |
| OmniGenome | SNT | |
| OmniGenome | SNT |
We include a popular bioinformatics baseline of RNAInverse and select several recent DNA and RNA FMs which support masked language modelling. We exclude HyenaDNA in this experiment because it does not support masked nucleotide prediction. It is observed from Table 3 that RNAInverse solved of the RNA design puzzles, indicating a promising capability in RNA design. The FMs, such as 3UTRBERT and DNABERT2 fail in RNA design because they cannot handle SN-level modelling. Meanwhile, RNA-MSM, RNA-FM and SpliceBERT demonstrated trivial proficiency in RNA design, solving to puzzles. This observation suggests these FMs cannot precisely predict the bases without any Str2Seq prediction ability. With the help of Str2Seq, i.e., structure-contextualised sequence reconstruction, OmniGenome and OmniGenome significantly outperformed other FMs with and puzzles solved, respectively, underscoring the significance of Str2Seq in sequence-structure alignment. Besides, we expect an increase in performance with sufficient computational budgets and the findings provide crucial evidence of the significance of Str2Seq for RNA sequence design.
Zero-shot SSP (Seq2Str) Evaluation
This subsection evaluates both Seq2Str and Str2Seq prediction in sequence-structure alignment. The evaluation of Seq2Str is based on zero-shot SSP. We use OmniGenome and OmniGenome without fine-tuning to predict the secondary structures of sequences from the testing datasets and measure the macro-F1 score, where better structure prediction performance indicates a stronger capability for Seq2Str prediction. The experimental results are available in Table 4.
| Model | Zero-shot SSP (F1) | ||
| Archive2 | bpRNA | Stralign | |
| ViennaRNA | |||
| OmniGenome | |||
| OmniGenome | |||
| OmniGenome | |||
| OmniGenome | |||
| Model | PolyA | LncRNA | Chrom Acc | Prom Str | Term Str | Splice | Gene Exp | Enhancer |
| F1 | F1 | F1 | RMSE | RMSE | F1 | RMSE | F1 | |
| DNABERT2 | ||||||||
| HyenaDNA | ||||||||
| Caduceus | ||||||||
| NT-V2 | ||||||||
| Agro-NT | ||||||||
| SpliceBERT | ||||||||
| 3UTRBERT | ||||||||
| RNA-BERT | ||||||||
| RNA-MSM | ||||||||
| RNA-FM | ||||||||
| OmniGenome | ||||||||
| OmniGenome | ||||||||
| OmniGenome | ||||||||
| OmniGenome |
The results in Table 4 indicate that OmniGenome FMs mirrored the zero-shot secondary structure prediction (i.e., Seq2Str) performance of ViennaRNA. Moreover, OmniGenome and OmniGenome outperform OmniGenome FMs based on structure contexts from ViennaRNA. Given the ablation of structure contexts, OmniGenome also achieves performance comparable with ViennaRNA on the Archive2, bpRNA and RNAStralign datasets. Besides, we found that OmniGenome generally obtains better performance on a wide genome downstream tasks owing to the structure awareness, and random or noise structure contexts have no obvious effects on the structure prediction. We cannot compare with other FMs in the zero-shot SSP experiments, because existing FMs were not pertained for secondary structure prediction.
3.4 Results on RGB
| Model | mRNA | SNMD | SNMR | Archive2 | Stralign | bpRNA |
| RMSE | AUC | F1 | F1 | F1 | F1 | |
| ViennaRNA | N.A. | N.A. | N.A. | |||
| MXFold2 | N.A. | N.A. | N.A. | |||
| Ufold | N.A. | N.A. | N.A. | |||
| DNABERT2 | ||||||
| HyenaDNA | ||||||
| Caduceus | ||||||
| NT-V2 | ||||||
| Agro-NT | ||||||
| SpliceBERT | ||||||
| 3UTRBERT | ||||||
| RNABERT | ||||||
| RNA-MSM | ||||||
| RNA-FM | ||||||
| OmniGenome | ||||||
| OmniGenome | ||||||
| OmniGenome | ||||||
| OmniGenome |
The results in Table 6 demonstrate the performance of OmniGenome and its generalizability across various fine-grained RNA downstream tasks. It is observed that OmniGenome models achieve better results than both RNA and DNA FM baselines, including Agro-NT and DNABERT2, which contain hundreds of millions of parameters. This is because the existing FMs usually adopt k-mers or BPE tokenization that cannot handle SN resolution tasks, e.g., single nucleotide mutation detection and repair, and structure prediction. Because of the Seq2Str pre-training, OmniGenome and OmniGenome models exhibit strong results in secondary structure prediction, underscoring OmniGenome’s capabilities in SN-level RNA sequence understanding and manipulation.
3.5 Results on PGB
The PGB is a plant-oriented genomic benchmark. Although the benchmark datasets in PGB are DNA-based tasks, we can still evaluate the performance of OmniGenome and its generalizability on multi-modal (i.e., DNA and RNA) genomic tasks. The results in Table 5 reveal substantial variability in the performance of different FMs, where OmniGenome outperformed other baseline models across most tasks, particularly in tasks like Polyadenylation, Splice Site, and Enhancer Region classification, where they achieved the highest F1 scores. This suggests that OmniGenome’s architecture is particularly adept at handling complex genomic sequences. In comparison, existing FMs, e.g., NT-V2 and Agro-NT, showed lower performance with more parameters than OmniGenome. Besides, the performance of OmniGenome suggests that the structure context can further enhance the performance of genomic modelling. Overall, OmniGenome models achieve state-of-the-art performance on both benchmarks, especially for OmniGenome variants. The results underscore the importance of sequence-structure alignment in achieving complex genomic modelling tasks.
4 Related Works
Biological sequence modelling, including DNA, RNA, and protein, has attracted attention in recent years. Protein modelling, e.g., AlphaFold [36, 37, 38] and ESM [39], has been studied for many years compared to DNA and RNA modelling. In the realm of genomic sequence modelling, several early works aimed at addressing diversified genome downstream subtasks. For instance, DNABERT [40] adapts the architecture of BERT [27] for genomic sequence modelling, showing preliminary performance for in-silico genomic tasks. DNABERT2 [22], a multi-species FM improved based on DNABERT, proposes replacing k-mers tokenization with BPE tokenization to improve model performance. To explore the performance of large-scale FMs, the nucleotide transformers (V1 & V2) [26], AgroNT [20] and SegmentNT [41] leveraged billions of parameters to boost genomic sequence modelling and achieved promising performance in understanding DNA genome, with model scales up to billion and billion parameters, respectively. Agro-NT [20] was pre-trained on multi-species edible plant DNA sequences but failed to transfer effectively to RNA sequence modelling in our experiments. To address the modelling capacity problem caused by the remarkable lengths of genomes, there is a growing focus on the necessity of long-range sequence modelling and the introduction of autoregressive FMs, namely, HyenaDNA [1] and Evo [42]. In terms of RNA genomic modelling, there have been some preliminary works, such as scBERT [43], RNABERT [44], RNA-FM [12], RNA-MSM [45], and RNAErnie [46], to name a few. However, these methods have only trained the FMs on a limited-scale database, as RNA sequences are generally expensive to obtain. Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) [47], 5’ untranslated regions (5’UTR) [35], 3’ untranslated regions (3’UTR) [25], or precursor mRNA sequences [24], thus limiting the models’ ability to capture the diversity of RNA sequences. Uni-RNA [34] has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM [35] adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training. However, these FMs are not available for Str2Seq mapping and suffer from limited model and data scales that fail to uncover the comprehensive efficacy of sequence-structure alignment on a wide set of genomic tasks. ERNIE-RNA [48] feeds the RNA structure along with the sequence into the model and improves the downstream tasks. However, it also ignores the significance of Str2Seq prediction capability. In a nutshell, existing FMs fail to achieve sequence-structure alignment without exception.
5 Conclusion
We introduced OmniGenome to tackle the challenge of sequence-structure alignment in genome modelling, which bridges the gap between sequence and structural information and improves the reliability of genome analysis. Experimental results on four comprehensive in-silico RNA and DNA benchmarks demonstrate that OmniGenome outperforms existing FMs across diversified downstream tasks, e.g., up to F1 score for SSP and accuracy of RNA design. The superior performance highlights the potential of sequence-structure alignment in the field of genomics.
Acknowledgment
[…]

### v3

@@
-OmniGenome: Aligning RNA Sequences with Secondary Structures in Genomic Foundation Models
-Abstract: The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated.
+Bridging Sequence-Structure Alignment in RNA Foundation Models
+Abstract
+The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated.
 Existing FMs have struggled to establish sequence-structure alignment, hindering the free flow of genomic information between RNA sequences and structures.
@@
 We leverage four comprehensive in-silico genome modelling benchmarks to evaluate performance across a diverse set of genome downstream tasks, where the results show that OmniGenome achieves state-of-the-art performance on RNA and DNA benchmarks, even without any training on DNA genomes.
-Keywords: RNA Foundation Model, RNA Design, Secondary Structure Prediction
 1 Introduction
@@
 To decipher the genomic code in RNA and manipulate RNA engineering and design, current research mainly uses bioinformatics in solving RNA genome-oriented challenges.
-Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ [1] hidden in genome encoding processes.
-Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction [2, 3, 4, 5], degradation rate prediction [6, 7], and mRNA vaccine design [8, 9].
-In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology [10].
+Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ (Nguyen et al.
+2023) hidden in genome encoding processes.
+Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction (Tan et al.
+2017; Danaee et al.
+2018; Mathews 2019; Kalvari et al.
+2021), degradation rate prediction (Yaish and Orenstein 2022; Wayment-Steele et al.
+2022), and mRNA vaccine design (Corbett et al.
+2020; Runge et al.
+2023).
+In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology (Ganser et al.
+2019).
 However, the role of the structure as a second ‘genomic language’ to interact with sequences and solve various RNA downstream tasks has been largely ignored.
-Sequence-Structure Alignment in FMs
+Sequence-Structure Alignment in GFMs
 We define alignment between sequences and secondary structures111In this paper, all structures are referred to as the secondary structures.
@@
 There has been a deep scientific challenge to align RNA sequences with structures because it is not deterministic to predict sequences from structures and vice versa.
-In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors [11].
-Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Fig.
-1.
+In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors (Tinoco Jr and Bustamante 1999).
+Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Figure 1.
 The oversight of such alignment in existing FMs causes outstanding issues in understanding and leveraging RNA structures, such as mRNA design.
-For example, recent state-of-the-art RNA FMs, RNA-FM [12] and RNA-MSM [13], only solved out of puzzles in in-silico RNA design [14].
+For example, recent state-of-the-art RNA FMs, RNA-FM (Chen et al.
+2022) and RNA-MSM (Zhang et al.
+2024), only solved out of puzzles in in-silico RNA design (Lee et al.
+2014).
 This is because they fail to decipher corresponding sequences based on structures to guide RNA design.
@@
 To induce the ability of Str2Seq mapping in genomic FMs, we formulate a structure-contextualised RNA sequence reconstruction task, which stems from the representation of RNA secondary structures in texts composed of dots and brackets.
-As the diagram is shown in Fig.
-2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence.
+As the diagram is shown in Figure 2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence.
 Then, we pre-train OmniGenome to reconstruct the masked nucleotide bases given the structure contexts.
@@
 Seq2Str Mapping
-On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) [15, 16], is another critical aspect of achieving the alignment.
-We generalise end-to-end structure pre-training [17] to OmniGenome pre-training.
+On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) (Sato, Akiyama, and Sakakibara 2021; Fu et al.
+2022), is another critical aspect of achieving the alignment.
+We generalise end-to-end structure pre-training (Yan, Hamilton, and Blanchette 2022) to OmniGenome pre-training.
 This large-scale structure pre-training on diversified genomes supervises OmniGenome to perform Seq2Str mapping.
-The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions [18] and barriers the structure prediction robustness on small datasets.
+The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions (Chen et al.
+2020) and barriers the structure prediction robustness on small datasets.
 To conduct Seq2Str mapping, tremendous secondary structure annotations are required to avoid data bias.
 A feasible solution to RNA structure pre-training is leveraging the plausible structures calculated based on the minimum free energy.
-In this paper, we leverage the popular ViennaRNA [19] to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
-Evaluation and Results
+In this paper, we leverage the popular ViennaRNA (Lorenz et al.
+2011) to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
+Evaluations and Results
 To validate the effectiveness of OmniGenome, we designed four large-scale genome benchmarks with diverse genomics tasks.
 The first one is the RNA genomics benchmark (RGB) compiled in the study, which contains diverse challenging genomics understanding tasks that benefit from the sequence-structure alignment, such as degradation rate prediction.
-The second benchmark is the plant genomics benchmark (PGB) [20] which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks.
+The second benchmark is the plant genomics benchmark (PGB) (Mendoza-Revilla et al.
+2023) which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks.
 In particular, we want to use this benchmark to evaluate the generalisability of OmniGenome among diversified species and genomes.
-The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT [20] which contains billion parameters.
-The last two benchmarks, available in the appendix, are the genomics benchmark (GB) [21] and genomics understanding evaluation (GUE) [22], which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
+The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT (Mendoza-Revilla et al.
+2023) which contains billion parameters.
+The last two benchmarks, available in the appendix, are the genomics benchmark (GB) (Grešová et al.
+2023) and genomics understanding evaluation (GUE) (Zhou et al.
+2023), which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
 In addition, we also conduct zero-shot Seq2Str and Str2Seq prediction experiments to verify the performance of sequence-structure alignment.
@@
 In terms of Str2Seq prediction performance, we evaluate the performance of OmniGenome in the in-silico RNA design task.
-We solved of complex puzzles of the EternaV2 benchmark [14], while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to .
+We solved of complex puzzles of the EternaV2 benchmark (Lee et al.
+2014), while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to .
 Besides, OmniGenome only takes less than one hour to solve most of the puzzles, while most RNA design methods need to take up to hours to solve even a single puzzle.
@@
 Open science is always the golden standard to promote this rising area of FM for genome modelling, which unfortunately lacks relevant high-quality resources such as code integrity, data availability, and pre-training pipeline.
-To address this gap, following the FAIR principles [23], we developed an open-source package222https://tinyurl.com/OmniGenome that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few.
+To address this gap, following the FAIR principles (Wilkinson et al.
+2016), we developed an open-source package222https://github.com/yangheng95/OmniGenBench that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few.
 It provides ready-to-use genomics benchmarks and uses the API with only a few lines of code to streamline benchmarking purposes.
@@
 We aim to implement a fine-grained alignment between RNA sequences and structures, where each base in the sequences reflects a structural label in ‘(’, ‘)’, ‘.’.
-Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method [1, 24] in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels.
-We illustrate our tokenization based on an example shown in Fig.
-3.
+Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method (Nguyen et al.
+2023; Chen et al.
+2023) in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels.
+We illustrate our tokenization based on an example shown in Figure 3.
 Our adapted SNT features bidirectional mappings between single nucleotide (SN) bases and structural labels required by sequence-structure alignment.
-Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution [1, 24].
-For example, the k-mers [25, 26] and BPE [27, 22] tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction.
-Like other encoder-only models, e..g, BERT [27], we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
+Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution (Nguyen et al.
+2023; Chen et al.
+2023).
+For example, the k-mers (Yang et al.
+2023; Dalla-Torre et al.
+2023) and BPE (Devlin et al.
+2019; Zhou et al.
+2023) tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction.
+Like other encoder-only models, e..g, BERT (Devlin et al.
+2019), we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
 2.2 Pre-training Objectives
@@
 OmniGenome adopts the Transformer encoder architecture with bidirectional multi-head attention.
-We do not adopt recent architectures like Mamba [28, 29] and Hyena [1] because our experiments in Table 5 and Table 6 show that these architectures are not competent at RNA genome understanding.
+We do not adopt recent architectures like Mamba (Gu and Dao 2023; Schiff et al.
+2024) and Hyena (Nguyen et al.
+2023) because our experiments in Table 4 and Table 5 show that these architectures are not competent at RNA genome understanding.
 This low performance is probably because RNA sequences are much shorter than DNA sequences in the wild.
@@
 The batch size is set to .
-No dropout is applied during pre-training, and we use the rotary position embeddings [30] to further enhance the model’s scalability to long RNA sequences.
-We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix 6.
+No dropout is applied during pre-training, and we use the rotary position embeddings (Su et al. 2024) to further enhance the model’s scalability to long RNA sequences.
+We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix C.
 The pre-training was finished in approximately 1 and 3 weeks for OmniGenome and OmniGenome, respectively.
 2.4 Pre-training Database: OneKP
-Recent studies [24, 22] have shown that data diversity can enhance FM performance without significantly increasing model capacity.
-For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ [31], which compiles large-scale RNA raw sequence database from plant species.
+Recent studies (Chen et al.
+2023; Zhou et al.
+2023) have shown that data diversity can enhance FM performance without significantly increasing model capacity.
+For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ (Carpenter, Leebens-Mack, and et al.
+2019), which compiles large-scale RNA raw sequence database from plant species.
 The raw sequences are not available for pre-training before processing and filtering.
@@
 These multi-species and SN-level tasks in RGB serve as the first comprehensive RNA benchmark to assess the modelling capabilities of FMs.
-For detailed information on each dataset, such as their sources and sizes, please refer to Appendix 9.1.
+For detailed information on each dataset, such as their sources and sizes, please refer to Appendix F.1.
 Plant Genomic Benchmark (PGB)
-PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark [20] ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology.
+PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark (Mendoza-Revilla et al.
+2023) ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology.
 PGB involves types of DNA downstream subtasks, including a range of critical tasks such as promoter strength prediction and gene expression regression.
@@
 In this section, we provide a case of RNA design that requires the exploitation of sequence-structure alignment.
-To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/OmniGenome/tutorials/zeroshot_omnigenome_RNA_design.py to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome.
-The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix 10 and Fig.
-5, respectively.
+To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/RNADesign to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome.
+The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix G and Figure 5, respectively.
 In Section 3.3, the experimental results on the Eterna V2 dataset indicate an impressive performance in RNA design compared to existing methods.
@@
 The e-value has been set to for rigorous sequence filtering.
-| Model | Tokenization | # of Params | Pre-training Data Scale | Pre-training Data Source | Species | Sequence Type |
-| DNABERT-2 | BPE | M | B Tokens | The 1000 Genomes Project | Human + Species | DNA |
-| NT-V2-M | k-mers | M | B Tokens | The 1000 Genomes Project, etc. | Human + Species | DNA |
-| HyenaDNA-Large | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
-| Caduceus | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
-| Agro-NT-B | k-mers | M | B Tokens | Ensembl Plants Database | Edible Plants | DNA |
-| SpliceBERT | SNT | M | M Sequences | UCSC Genome Browser | Multi-Vertebrates | precursor-mRNA |
-| RNA-BERT | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
-| RNA-MSM | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
-| RNA-FM | SNT | M | M Sequences | RNAcentral Database | Multi-Species | ncRNA |
-| 3UTRBERT | k-mers | M | Sequences | The GENCODE Project | Human | mRNA ’UTR |
-| OmniGenome | SNT | M | B Tokens | The OneKP Initiative | Plant Species | mRNA, CDS, UTR |
-| OmniGenome | M |
 3.2 Comparison Baselines
@@
 In SSP tasks, we can also use the ViennaRNA’s structure annotations as contexts to improve downstream SSP performance.
-Please refer to Appendix 8 for brief introductions of these FMs.
+Please refer to Appendix E for brief introductions of these FMs.
 We can compare OmniGenome with the following RNA and DNA FMs shown in Table 6 as baselines to help evaluate the performance of OmniGenome.
-We are aware that some FMs are also developed for RNA, such as Uni-RNA [34], 5UTR-LM [35], etc.
+We are aware that some FMs are also developed for RNA, such as Uni-RNA (Wang et al.
+2023), 5UTR-LM (Chu et al.
+2024), etc.
 However, we cannot compare OmniGenome with them because their source codes are very hard to work with in our efforts or not publicly available.
@@
 we demonstrate the Str2Seq prediction capability of OmniGenome based on RNA design.
-We employed the Eterna777https://github.com/eternagame/eterna100-benchmarking/blob/master/eterna100_vienna2.txt [14] V2 benchmark, which consists of specified secondary structures.
+We employed the Eterna (Lee et al.
+2014) V2 benchmark, which consists of specified secondary structures.
 This task aims to design RNA sequences based on reference structures.
 We develop a genetic algorithm (GA) which exploits masked nucleotide modelling (a.k.a., masked language modelling) to find plausible RNA sequences that solve RNA design puzzles.
-The implementation details can be found in Fig.
-5 in Appendix 10.
+The implementation details can be found in Figure 5 in Appendix G.
 In the GA, the population size is set at , with iterations, and the mutation rate for each base is .
@@
 | OmniGenome | SNT |
-We include a popular bioinformatics baseline of RNAInverse and select several recent DNA and RNA FMs which support masked language modelling.
+We include a popular baseline of RNAInverse and select recent DNA and RNA FMs which support masked language modelling.
 We exclude HyenaDNA in this experiment because it does not support masked nucleotide prediction.
@@
 4 Related Works
-Biological sequence modelling, including DNA, RNA, and protein, has attracted attention in recent years.
-Protein modelling, e.g., AlphaFold [36, 37, 38] and ESM [39], has been studied for many years compared to DNA and RNA modelling.
-In the realm of genomic sequence modelling, several early works aimed at addressing diversified genome downstream subtasks.
-For instance, DNABERT [40] adapts the architecture of BERT [27] for genomic sequence modelling, showing preliminary performance for in-silico genomic tasks.
-DNABERT2 [22], a multi-species FM improved based on DNABERT, proposes replacing k-mers tokenization with BPE tokenization to improve model performance.
-To explore the performance of large-scale FMs, the nucleotide transformers (V1 & V2) [26], AgroNT [20] and SegmentNT [41] leveraged billions of parameters to boost genomic sequence modelling and achieved promising performance in understanding DNA genome, with model scales up to billion and billion parameters, respectively.
-Agro-NT [20] was pre-trained on multi-species edible plant DNA sequences but failed to transfer effectively to RNA sequence modelling in our experiments.
-To address the modelling capacity problem caused by the remarkable lengths of genomes, there is a growing focus on the necessity of long-range sequence modelling and the introduction of autoregressive FMs, namely, HyenaDNA [1] and Evo [42].
-In terms of RNA genomic modelling, there have been some preliminary works, such as scBERT [43], RNABERT [44], RNA-FM [12], RNA-MSM [45], and RNAErnie [46], to name a few.
+Current RNA FMs focused on sequence-to-structure mapping, e.g., end-to-end secondary structure prediction.
+However, to the best of our knowledge, the sequence-structure alignment in RNA genome modelling has yet been investigated in the literature.
+There have been some preliminary works, such as scBERT (Yang et al.
+2022), RNABERT (Akiyama and Sakakibara 2022), RNA-FM (Chen et al.
+2022), RNA-MSM (Zhang et al.
+2023), and RNAErnie (Wang et al.
+2024), to name a few.
 However, these methods have only trained the FMs on a limited-scale database, as RNA sequences are generally expensive to obtain.
-Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) [47], 5’ untranslated regions (5’UTR) [35], 3’ untranslated regions (3’UTR) [25], or precursor mRNA sequences [24], thus limiting the models’ ability to capture the diversity of RNA sequences.
-Uni-RNA [34] has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
-In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM [35] adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training.
+Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) (Hallee, Rafailidis, and Gleghorn 2023), 5’ untranslated regions (5’UTR) (Chu et al.
+2024), 3’ untranslated regions (3’UTR) (Yang et al.
+2023), or precursor mRNA sequences (Chen et al.
+2023), thus limiting the models’ ability to capture the diversity of RNA sequences.
+Uni-RNA (Wang et al.
+2023) has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
+In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM (Chu et al.
+2024) adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training.
 However, these FMs are not available for Str2Seq mapping and suffer from limited model and data scales that fail to uncover the comprehensive efficacy of sequence-structure alignment on a wide set of genomic tasks.
-ERNIE-RNA [48] feeds the RNA structure along with the sequence into the model and improves the downstream tasks.
+ERNIE-RNA (Yin et al.
+2024) feeds the RNA structure along with the sequence into the model and improves the downstream tasks.
 However, it also ignores the significance of Str2Seq prediction capability.
@@
 The superior performance highlights the potential of sequence-structure alignment in the field of genomics.
-Acknowledgment
+Acknowledgements
 […]
+Appendix A Reproducibility Checklist
+This paper:
+Includes a conceptual outline and/or pseudocode description of AI methods introduced yes
+Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results yes
+Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper yes
+Does this paper make theoretical contributions? yes
+If yes, please complete the list below.
+All assumptions and restrictions are stated clearly and formally. yes
+All novel claims are stated formally (e.g., in theorem statements). yes
+Proofs of all novel claims are included. yes
+Proof sketches or intuitions are given for complex and/or novel results. yes
+Appropriate citations to theoretical tools used are given. yes
+All theoretical claims are demonstrated empirically to hold. yes
+All experimental code used to eliminate or disprove claims is included. yes
+Does this paper rely on one or more datasets? yes
+If yes, please complete the list below.
+A motivation is given for why the experiments are conducted on the selected datasets yes
+All novel datasets introduced in this paper are included in a data appendix. yes
+All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. yes
+All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. yes
+All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. yes
+All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. yes
+Does this paper include computational experiments? yes
+If yes, please complete the list below.
+Any code required for pre-processing data is included in the appendix. yes
+All source code required for conducting and analyzing the experiments is included in a code appendix. yes
+All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. yes
+All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from yes
+If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. yes
+This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks.
+yes
+This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. yes
+This paper states the number of algorithm runs used to compute each reported result. yes
+Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. yes
+The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). N.A. because of computation complexity
+This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. yes
+This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. yes
+Appendix B DNA Foundation Models
+Biological sequence modelling, including DNA, RNA, and protein, has attracted attention in recent years.
+Protein modelling, e.g., AlphaFold (Jumper et al.
+2021; Evans et al.
+2021; Abramson et al.
+2024) and ESM (Lin et al.
+2022), has been studied for many years compared to DNA and RNA modelling.
+In the realm of genomic sequence modelling, several early works aimed at addressing diversified genome downstream subtasks.
+For instance, DNABERT (Ji et al.
+2021) adapts the architecture of BERT (Devlin et al.
+2019) for genomic sequence modelling, showing preliminary performance for in-silico genomic tasks.
+DNABERT2 (Zhou et al.
+2023), a multi-species FM improved based on DNABERT, proposes replacing k-mers tokenization with BPE tokenization to improve model performance.
+To explore the performance of large-scale FMs, the nucleotide transformers (V1 & V2) (Dalla-Torre et al.
+2023), AgroNT (Mendoza-Revilla et al.
+2023) and SegmentNT (de Almeida et al.
+2024) leveraged billions of parameters to boost genomic sequence modelling and achieved promising performance in understanding DNA genome, with model scales up to billion and billion parameters, respectively.
+Agro-NT (Mendoza-Revilla et al.
+2023) was pre-trained on multi-species edible plant DNA sequences but failed to transfer effectively to RNA sequence modelling in our experiments.
+To address the modelling capacity problem caused by the remarkable lengths of genomes, there is a growing focus on the necessity of long-range sequence modelling and the introduction of autoregressive FMs, namely, HyenaDNA (Nguyen et al.
+2023) and Evo (Nguyen et al.
+2024).
+Appendix C Pre-training Environment
+The pre-training of OmniGenome was conducted on a dedicated Linux computation node, equipped with Nvidia RTX GPUs.
+For distributed model training, we employed version of the Transformers library alongside version of the Accelerate library.
+Our implementation framework of choice for OmniGenome was PyTorch, specifically version .
+The ViennaRNA version is in our experiments.
+While some existing code was adapted for the modules within OmniGenome, the majority of the codebase, such as genomic sequences preprocessing, model pre-training, objective functions, and experiments, was meticulously crafted from scratch.
+Appendix D OmniGenome Package
+Genome modelling is still in its early stages.
+Therefore, open-source codes and resources are consequently very scarce.
+Typically, existing FMs usually release the pre-trained model weights only, without providing the training and fine-tuning codes, and benchmark evaluation scripts, etc.
+To address this issue, we have developed a comprehensive open-source RNA genome modelling toolkit777https://github.com/yangheng95/OmniGenBench based on OmniGenome.
+This toolkit aims to provide extensive FM fine-tuning tutorials and a unified automated benchmark evaluation.
+The main features of the OmniGenome Package are as follows:
+Fine-Tuning Tutorials: We provide tutorials for fine-tuning on all downstream genome modelling tasks, including dataset processing, model implementation, and training processes.
+A fine-tuning example for secondary structure is included, covering both training and demonstration of secondary structure prediction.
+Automated Benchmark Evaluation: We offer an automated benchmark evaluation interface, which includes the built-in PGB and RGB benchmarks.
+By predefining the configurations for benchmark evaluation subtasks, such as hyperparameters, our tool supports the automated benchmark evaluation of future FMs and the addition of new benchmarks.
+The goal of automated benchmark evaluation is to ensure fairness and ease of use.
+We provide a tutorial on automated evaluation to guide users in benchmark evaluation.
+The automated benchmarking tutorial is available at: https://github.com/yangheng95/OmniGenBench/blob/master/examples/benchmark/autobench.ipynb.
+Online Hub for Genome Modelling: We have created a hub for hosting and distributing open-source licensed datasets, model checkpoints, and benchmark evaluations.
+Additionally, we have designed flexible interfaces to support the sharing of datasets and models within the community.
+This approach helps mitigate the issue of resource scarcity.
+The hub will be available soon.
+We are in the process of finalising the necessary documentation and will officially release this tool in the near future.
+| Model | Tokenization | # of Params | Pre-training Data Scale | Pre-training Data Source | Species | Sequence Type |
+| DNABERT-2 | BPE | M | B Tokens | The 1000 Genomes Project | Human + Species | DNA |
+| NT-V2-M | k-mers | M | B Tokens | The 1000 Genomes Project, etc. | Human + Species | DNA |
+| HyenaDNA-Large | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
+| Caduceus | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
+| Agro-NT-B | k-mers | M | B Tokens | Ensembl Plants Database | Edible Plants | DNA |
+| SpliceBERT | SNT | M | M Sequences | UCSC Genome Browser | Multi-Vertebrates | precursor-mRNA |
+| RNA-BERT | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
+| RNA-MSM | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
+| RNA-FM | SNT | M | M Sequences | RNAcentral Database | Multi-Species | ncRNA |
+| 3UTRBERT | k-mers | M | Sequences | The GENCODE Project | Human | mRNA ’UTR |
+| OmniGenome | SNT | M | B Tokens | The OneKP Initiative | Plant Species | mRNA, CDS, UTR |
+| OmniGenome | M |
+Appendix E Comparison Baselines
+There are no direct counterparts to OmniGenome aimed at plant RNA genome modelling.
+We can compare it with the following recent RNA and DNA genome FMs as potential baselines to help evaluate the performance of OmniGenome.
+The brief introductions of the FMs in Table 6 are as follows:
+ViennaRNA (Lorenz et al.
+2011).
+ViennaRNA is a comprehensive genomic analysis tool that includes a diverse set of interfaces, such as RNAFold888https://www.tbi.univie.ac.at/RNA/RNAfold.1.html and RNAInverse999https://www.tbi.univie.ac.at/RNA/RNAinverse.1.html design.
+ViennaRNA serves as the baseline for RNA structure prediction and RNA design in our experiments.
+MXFold2 (Sato, Akiyama, and Sakakibara 2021).
+MXFold2 is a deep learning model for RNA secondary structure prediction, integrating thermodynamic models to improve prediction accuracy.
+It stands out for its robustness across various RNA families and is widely used for predicting RNA secondary structures from given sequences.
+UFold (Fu et al.
+2022).
+UFold is a deep learning-based RNA secondary structure prediction tool designed for speed and accuracy.
+It uses a convolutional neural network (CNN) architecture optimized to predict base-pairing probabilities in RNA sequences, enabling efficient secondary structure prediction.
+DNABERT2 (Zhou et al.
+2023).
+DNABERT2 is one of the latest DNA FMs which improves the performance of DNABERT.
+The main modification of DNABERT2 is the tokenization method, which was changed to BPE from k-mers.
+HyenaDNA (Nguyen et al.
+2023).
+HyenaDNA is an autoregressive FM optimised for long-range genome data processing.
+HyenaDNA is based on the Hyena convolution architecture and capable of handling sequences up to M bases in length.
+Caduceus (Schiff et al.
+2024).
+Caduceus101010https://huggingface.co/kuleshov-group/caduceus-ps˙seqlen-131k˙d˙model-256˙n˙layer-16 is an advanced DNA language model built on the MambaDNA architecture, designed to address challenges in genomic sequence modelling, such as long-range token interactions and reverse complementarity (RC).
+Nucleotide Transformer (NT) V2 (Dalla-Torre et al.
+2023).
+The NT FMs were trained on DNA data, including the human reference genome and multi-species DNA sequences.
+They aim to capture the complex patterns within nucleotide sequences for various genome modelling applications.
+Agricultural Nucleotide Transformer (Agro-NT) (Mendoza-Revilla et al. 2023). Agro-NT is a large-scale DNA FM (B parameters) akin to the Nucleotide Transformers but with a focus on plant DNA.
+SpliceBERT (Chen et al. 2023). It was trained on M precursor messenger RNA (pre-mRNA) and specialized in RNA splicing of pre-mRNA sequences.
+3UTRBERT (Yang et al.
+2023).
+This model was trained on k 3’UTRs for 3’UTR-mediated gene regulation tasks.
+It uses k-mers tokenization instead of SNT.
+RNA-BERT (Akiyama and Sakakibara 2022).
+RNA-BERT is a BERT-style model pre-trained on a large corpus of non-coding RNA sequences.
+It uses masked language modelling (MLM) as its primary training objective.
+The model is designed to predict RNA structural alignments and can be fine-tuned for various RNA sequence classification and regression tasks
+RNA-MSM (Zhang et al.
+2024) RNA-MSM is an unsupervised RNA language model based on multiple sequence alignment (MSA).
+It is the first model of its kind to produce embeddings and attention maps that directly correlate with RNA secondary structure and solvent accessibility.
+RNA-MSM is particularly effective for tasks involving evolutionary relationships in RNA sequences.
+RNA-FM (Chen et al.
+2022) RNA-FM is a BERT-based RNA foundation model trained on a vast dataset of non-coding RNA sequences.
+The model excels in predicting RNA structure and function by leveraging masked language modelling (MLM) during pre-training.
+RNA-FM’s training data is sourced from the RNAcentral database, providing it with extensive knowledge across diverse RNA species.
+OmniGenome. OmniGenome is the RNA genome FM that advocates the importance of sequence-structure alignment. Moreover, it is the first FM which addressed the in-silico RNA design task.
+OmniGenome121212Please find the finetuning example111111https://anonymous.4open.science/r/OmniGenome-SupplementaryMaterials-002F/tutorials/ssp˙finetuning.py of OmniGenome in the supplemental materials..
+OmniGenome is a variant of OmniGenome that feeds both sequences and structures into OmniGenome to aggregate the feature representations to improve modelling ability.
+These genomic FMs serve as valuable baselines for evaluating the performance of OmniGenome in our study, particularly in RNA structure prediction and sequence-to-structure mapping tasks.
+Appendix F Benchmark Suites
+In this section, we first introduce the details, such as downstream task types and species, of RGB and PGB.
+Besides, we have included two recent DNA benchmarks in the evaluation, aiming to provide comprehensive performance comparisons between OmniGenome with existing DNA FMs, where OmniGenome was not pre-trained on any specific DNA genome database.
+F.1 RNA Genomic Benchmark (RGB)
+RGB contains SN-level tasks that are curated or collected from the literature, and the details of the RGB can be found in Table 7.
+The objective of RGB is to benchmark genome FMs in challenging SN-level modelling tasks such as the detection and repair of SN mutations, mRNA sequence degradation rates, and RNA secondary structure prediction.
+Due to the lack of a plant RNA benchmark dataset, RGB enriches the benchmark suites by including the modelling of RNA downstream tasks from a variety of species, e.g., plants and animals.
+The sequence length in RGB ranges from to , which is sufficient for most RNA understanding tasks.
+In summary, these multi-species and SN-level tasks in RGB serve as the first comprehensive benchmark utilised to assess the RNA sequence modelling capabilities of OmniGenome and its baseline models.
+The brief introduction of the datasets in RGB is as follows:
+Single-Nucleotide Mutation Detection (SNMD): We developed a plant RNA dataset synthesising the single-nucleotide mutations.
+Focused on identifying potential single nucleotide changes, this task is essential for detecting mutations linked to genetic disorders.
+The SNMD dataset introduces up to random mutations in the original sequences, regardless of variation ratios.
+Cross-entropy is utilised as the loss function for this binary token classification task.
+Single-Nucleotide Mutation Repair (SNMR): This task challenges the model to suggest corrective actions at the single nucleotide level, aiding in gene therapy approaches.
+The SNMR dataset mirrors the SNMD dataset, with cross-entropy as the loss function, indicating a token 4-way (i.e., A, U, C, G) classification task.
+mRNA Degrade Rate Prediction (mRNA): Estimating the decay rate of nucleotide bases in mRNA sequences, this task is vital for deciphering gene expression and regulation.
+The dataset originates from the Kaggle COVID-19 vaccine design competition131313https://www.kaggle.com/competitions/stanford-covid-vaccine, focusing solely on sequence-based degradation rate prediction and excluding RNA structures.
+It’s a token regression task using MSE as the loss function, with the dataset re-split into training, validation, and testing sets for evaluation.
+RNA Secondary Structure Prediction (bpRNA & Archive2 & RNAStralign): Aiming to predict RNA folding into secondary structures, this task is fundamental to RNA functionality and interactions.
+We evaluated OmniGenome on four datasets, bpRNA (Danaee et al.
+2018), ArchiveII (Mathews 2019), RNAStralign (Tan et al.
+2017) and Rfam (Kalvari et al.
+2021).
+Following existing works, we have excluded sequences over bases and complex structures, simplifying to three symbols: ‘(’, ‘.’, ‘)’Ẇe have filtered the RNAStralign, Archive2 and bpRNA datasets using CD-HIT-EST and BLAST as described in Section 3.1, which results in different data splits compared to existing works for SSP.
+Please find the dataset details of splits in Table 7.
+Besides, the SSP datasets processed in different publications are usually unknown because of various data filtering implementations.
+As a result, the performance of OmniGenome may not be directly compared with other studies.
+RNA SSP tasks are trained based on cross-entropy loss functions.
+| Task | Task Type | # of examples | # of classes | Metric | Sequence length | Source |
+| SNMD | Token classification | AUC | This work | |||
+| SNMR | Token classification | F1 | This work | |||
+| mRNA | Token regression | — | RMSE | Kaggle141414https://www.kaggle.com/competitions/stanford-covid-vaccine | ||
+| bpRNA | Token classification | F1 | (Franke et al. 2024) | |||
+| AchiveII | Token classification | F1 | (Mathews 2019) | |||
+| StrAlign | Token classification | F1 | (Tan et al. 2017) |
+Please find the appendix for the input and output examples of each subtask in RGB.
+The detailed task descriptions for each nucleic acid and species, including the number of examples, classes, evaluation metric, and sequence length, are outlined in Table 7.
+Each task is carefully curated to reflect the complexity and variety inherent in genomic data, providing a robust framework for assessing the nuanced capabilities of state-of-the-art RNA FMs.
+Table 8 show the virtual examples of different datasets in RGB. Please refer to our supplementary materials to find the datasets for more details.
+| Genome Type | Dataset | Examples | |
+| RNA | SNMD | Input Sequence | G A G T A T T G A G |
+| True Label | 0 0 1 0 0 0 0 1 0 0 | ||
+| Prediction | 0 0 0 0 0 0 0 1 0 0 | ||
+| SNMR | Input Sequence | T A C G A C T G A T | |
+| True Label | T A C A A G T A A T | ||
+| Prediction | T A C A A C T G A T | ||
+| mRNA | Input Sequence | G G A C | |
+| True Label | [0.1,0.3,0.2] [0.8,0.4,0.1][0.9,0.4,0.3] [0.5,0.2,0.6] | ||
+| Prediction | [0.1,0.3,0.2] [0.8,0.4,0.1][0.9,0.4,0.3] [0.5,0.2,0.6] | ||
+| bpRNA | Input Sequence | G G C G A C U U U U | |
+| True Label | ( ( ( ) ) ) | ||
+| Prediction | ( ( ( ( ) ) ) ) | ||
+| Archive2 | Input Sequence | A G U A G U U U G C U | |
+| True Label | ( ( ( ) ) ) | ||
+| Prediction | ( ( ( ) ) ) | ||
+| RNAStralign | Input Sequence | A G U A G U U U G C U | |
+| True Label | ( ( ( ) ) ) | ||
+| Prediction | ( ( ( ) ) ) | ||
+| Rfam | Input Sequence | A G U A G U U U G C U | |
+| True Label | ( ( ( ) ) ) | ||
+| Prediction | ( ( ( ) ) ) |
+F.2 Plant Genomic Benchmark (PGB)
+The Plant Genomic Benchmark (Mendoza-Revilla et al.
+2023) (PGB) provides a comprehensive suite of DNA downstream datasets designed to evaluate and improve the predictive capabilities of genomic FMs in plant biology.
+This benchmark, as shown in Table 9, encompasses a range of critical genomic tasks151515https://huggingface.co/InstaDeepAI/agro-nucleotide-transformer-1b, including binary classification, single and multi-variable regression, and multi-label classification, addressing various aspects of plant genomics such as RNA processing, gene expression, and chromatin accessibility.
+By integrating diverse genomic tasks, the PGB aims to facilitate advanced research and development in plant genomics, offering a robust platform for the assessment and enhancement of model performance across different plant species.
+To obtain a detailed description of PGB, please refer to Agro-NT (Mendoza-Revilla et al.
+2023).
+| Task | # of datasets | Task Type | Total # of examples | # of classes | Metric | Sequence length |
+| Polyadenylation | Sequence classification | F1 | ||||
+| Splice site | Sequence classification | F1 | ||||
+| LncRNA | Sequence classification | F1 | ||||
+| Promoter strength | Sequence regression | — | RMSE | |||
+| Terminator strength | Sequence regression | — | RMSE | |||
+| Chromatin accessibility | Multi-label classification | F1 | ||||
+| Gene expression | Multi-variable regression | — | RMSE | |||
+| Enhancer region | Sequence classification | F1 |
+F.3 Genomic Benchmarks
+The genomic benchmark (GB) is also a DNA-oriented FM benchmark suite, which can be used for generalisability evaluation of OmniGenome.
+It contains a well-curated collection of datasets designed for the classification of genomic sequences, focusing on regulatory elements across multiple model organisms.
+This collection facilitates robust comparative analysis and development of genomic FMs.
+The task names in the original repository are complex, we abbreviate the names as follows:
+DEM corresponds to “Demo Coding vs Intergenomic Seqs”
+DOW is for “Demo Human or Worm”
+DRE represents “Drosophila Enhancers Stark”
+HCE is short for “Human Enhancers Cohn”
+HEE denotes “Human Enhancers Ensembl”
+HRE abbreviates “Human Ensembl Regulatory”
+HNP shortens “Human Nontata Promoters”
+HOR is an abbreviation for “Human Ocr Ensembl”
+DME simplifies “Dummy Mouse Enhancers Ensembl”
+The brief statistics for each dataset included in the GUE benchmark are displayed in Table 12.
+Similar to GUE, we run the evaluation on a subset of GB, where for each task we randomly select at most k samples from the original splits, e.g., training, testing and validation (if any) sets.
+| Task | # of Sequences | # of Classes | Class Ratio | Median Length | Standard Deviation |
+| DME | |||||
+| DEM | |||||
+| DOW | |||||
+| DRE | |||||
+| HCE | |||||
+| HEE | |||||
+| HRE | |||||
+| HNP | |||||
+| HOR |
+| Model | DEM | DOW | DRE | DME | HCE | HEE | HRE | HNP | HOR |
+| F1 | F1 | F1 | F1 | F1 | F1 | F1 | F1 | F1 | |
+| DNABERT-2 | |||||||||
+| HyenaDNA | |||||||||
+| Caduceus | |||||||||
+| NT-V2 | |||||||||
+| SpliceBERT | |||||||||
+| 3UTRBERT | |||||||||
+| OmniGenome |
+The experimental results presented in Table 11 demonstrate that OmniGenome consistently achieves competitive performance across a diverse array of genomic tasks.
+Notably, OmniGenome excels in the Human Ensembl Regulatory (HRE) task with an F1 score of , outperforming other models like DNABERT-2 and HyenaDNA in this specific benchmark.
+Additionally, OmniGenome shows robust results in tasks involving enhancer predictions (HEE) and non-TATA promoters (HNP), underscoring its versatility and effectiveness in processing complex genomic sequences.
+These findings highlight the advanced capabilities of OmniGenome in handling intricate genomic data, contributing significantly to the field of genomic research.
+F.4 Genomic Understanding Evaluation
+The Genome Understanding Evaluation (Zhou et al.
+2023) serves as a DNA genomic benchmark, encompassing datasets across nine crucial genome analysis tasks applicable to a variety of species.
+Similar to PGB and GB, it is used for evaluating the generalisability of OmniGenome on DNA genome benchmarking.
+To thoroughly assess the capabilities of genome foundation models across sequences of varying lengths, tasks have been chosen with input lengths spanning from to .
+The brief statistics for each dataset included in the GUE benchmark are displayed in Table 12, and the task descriptions are available in Zhang et al.
+(2023).
+Due to resource limitations, we do not include large-scale FMs in this benchmark, e.g., agro-NT and CDSBERT.
+Besides, we run the evaluation on a subset of GUE, where for each task we randomly select at most 10k samples from the original splits, e.g., training, testing and validation (if any) sets.
+| Task | Metric | Datasets | Training | Validation | Testing |
+| Core Promoter Detection | macro F1 | tata | |||
+| notata | |||||
+| all | |||||
+| Promoter Detection | macro F1 | tata | |||
+| notata | |||||
+| all | |||||
+| Transcription Factor Prediction (Human) | macro F1 | wgEncodeEH000552 | |||
+| wgEncodeEH000606 | |||||
+| wgEncodeEH001546 | |||||
+| wgEncodeEH001776 | |||||
+| wgEncodeEH002829 | |||||
+| Splice Site Prediction | macro F1 | reconstructed | |||
+[…]
+[…]
+[…]
+| MelMafkDm2p5dStd | |||||
+[…]
+| Epigenetic Marks Prediction | macro F1 | H3 | |||
+| H3K14ac | |||||
+| H3K36me3 | |||||
+| H3K4me1 | |||||
+| H3K4me2 | |||||
+| H3K4me3 | |||||
+| H3K79me3 | |||||
+| H3K9ac | |||||
+| H4 | |||||
+| H4ac | |||||
+| Covid Variant Classification | macro F1 | Covid | |||
+| Enhancer Promoter Interaction | macro F1 | GM12878 | |||
+| HeLa-S3 | |||||
+| HUVEC | |||||
+| IMR90 | |||||
+| K562 | |||||
+| NHEK | |||||
+| Species Classification | macro F1 | fungi | |||
+| virus |
+The benchmark results on GUE can be found in Table 13.
+While OmniGenome does not consistently outperform other models across all datasets, it consistently demonstrates top-tier performance despite not being pre-trained on any DNA genome database.
+These results indicate that, although some FMs are optimised for specific genomic tasks (such as SpliceBERT for splice site detection), OmniGenome, which is specifically designed for RNA genomes, shows robust and versatile performance across a variety of tasks.
+The varying performance across different tasks and species suggests that genomic tasks could benefit from strong generalisability, provided that biological domain knowledge is incorporated into the training of FMs.
+| Model Performance (macro F1 Score) | |||||||
+| Model | Yeast EMP | Mouse TF-M | Virus CVC | Human TF-H | Human PD | Human CPD | Human SSP |
+| DNABERT-2 | |||||||
+| HyenaDNA | |||||||
+| Caduceus | |||||||
+| NT-V2 | |||||||
+| SpliceBERT | |||||||
+| 3UTRBERT | |||||||
+| OmniGenome | |||||||
+Appendix G Str2Seq Modelling Case: RNA Design
+G.1 Genetic Algorithm
+The working mechanism of our designed genetic algorithm based on OmniGenome is implemented as the following five-step process:
+Step . Given the target RNA secondary structure, we use OmniGenome to generate a set of candidate sequences .
+Step . If the termination criterion is not met, go to Step ; otherwise, output the current best sequence .
+Step . Based on , use single-point crossover and mutation to generate a population of offspring sequences .
+Step . Combine and to obtain , and use OmniGenome to predict the corresponding secondary structures of each sequence in . Evaluate the fitness values of sequences in .
+Step . Sort according to the fitness values and preserve the best sequences to constitute a new . Return to Step .
+Note that the fitness value of a sequence , denoted as , is evaluated as the Hamming distance of the RNA secondary structure predicted by OmniGenome against the target structure.
+The above genetic algorithm is not terminated until the sequence for the target RNA secondary structure is identified or the allocated computational budget is exhausted.
+As demonstrated in the zero-shot experiments in Table 2, OmniGenome models achieve top-tier performance, i.e., OmniGenome solved out of puzzles.
+We show several complex examples of puzzles from the EternaV2 design benchmark.
+According to Figure 6, puzzles #5 and #11 with approximately + bases are solved, while these puzzles are challenging to existing FMs.
+Even for puzzles that are not completely solved, e.g., puzzles #3 and #27, OmniGenome generates very similar structures, where the nucleotide base difference ratio between the designed structure and the target structure is only .
+This finding indicates the proficiency of OmniGenome models in solving challenging single-nucleotide resolution genome tasks.
+G.2 OmniGenome for Long-context RNA Modelling
+We study RNA genome modelling in this work, and most of the RNA sequences in the wild are short.
+However, in the case of evaluating the modelling performance on long-context RNA genomes, such as RNA sequences that contain more than nucleotide bases, we conduct an experiment to verify the SSP performance of OmniGenome for long sequences.
+To be more specific, we collected the bpRNA dataset (Danaee et al.
+2018) from the official website161616https://bprna.cgrb.oregonstate.edu and filtered it using the same protocol as the RGB SSP tasks.
+We re-split this bpRNA dataset based on a length threshold of bases; i.e., we used the sequences bases to form the testing set, and the rest of the sequences were split into training and validation sets.
+This variant of the bpRNA dataset is called bpRNA-L, and the results are available in Table 14.
+| Dataset | ViennaRNA | RNA-BERT | RNA-MSM | RNA-FM | OmniGenome |
+| bpRNA-L |
+The results indicate that OmniGenome outperforms existing RNA FMs in long-context RNA structure prediction.
+Specifically, OmniGenome achieved the highest F1 score of , surpassing other models such as ViennaRNA and RNA-FM.
+This demonstrates the capability of OmniGenome in handling complex RNA structures in sequences longer than bases, reinforcing the importance of sequence-structure alignment in RNA genome modelling.
+Appendix H Pre-training Objective Ablation Experiment
+We have included ablation experiments of pre-training objectives to study their effectiveness.
+For example, we ablated the Str2Seq and Seq2Str objectives for the OmniGenome variant.
+Due to resource considerations, we only trained each ablation with steps and evaluated the performance on RGB.
+The results are in Table 15.
+| OmniGenome-52M | mRNA | SNMD | SNMR | Archive2 | Stralign | bpRNA |
+| Ablations | RMSE | AUC | F1 | F1 | F1 | F1 |
+| MLM | ||||||
+| MLM+Str2Seq | ||||||
+| MLM+Seq2str | ||||||
+| MLM+Str2Seq+Seq2str |
+Overall, the results of the ablation experiments show that combining the masked language modelling (MLM) objective with both Str2Seq and Seq2Str mappings leads to the best performance across all evaluation metrics on the RGB benchmark.
+The model variant trained with all three objectives (Str2Seq, Seq2Str, and MLM) achieved the lowest RMSE and the highest AUC and F1 scores, indicating that these objectives are complementary and enhance the model’s ability to generalise to different RNA genomic tasks.
+Appendix I The OneKP Initiative
+There has been a variety of FMs utilised in different species, e.g., humans (Nguyen et al.
+2023; Dalla-Torre et al.
+2023), bacteria (Nguyen et al.
+2024), and viruses (Peng et al.
+2024), which indicates the effectiveness of pre-trained FMs on multi-species genomics.
+In this work, we aim to propose an FM for multi-species plant RNA sequence modelling.
+We leverage the OneKP initiative (Carpenter, Leebens-Mack, and et al.
+2019) to address the scarcity of plant RNA data, which contains species of plant transcriptomes.
+The scale of OneKP enables the development of a more robust and transferable RNA FM.
+The Plant Transcriptomes Initiative (OneKP) was a comprehensive effort aimed at exploring genetic diversity across the green plant kingdom (Viridiplantae), sequencing the RNA from ( in other versions) samples that represent over species, encompassing all major taxa within Viridiplantae.
+This includes streptophyte and chlorophyte green algae, bryophytes, ferns, angiosperms, and gymnosperms.
+The initiative’s final or capstone publication presents three major analyses: inferring species trees, identifying whole genome duplications, and detecting gene family expansions.
+These findings are particularly valuable for plant and evolutionary scientists interested in specific gene families, whether their focus is across the entire green plant tree of life or within more narrowly defined lineages.
+The sampling strategy for the 1KP was global and collaborative, with samples sourced from a wide range of environments including wild field collections, greenhouses, botanical gardens, laboratory specimens, and algal culture collections.
+The initiative prioritised the collection of live growing cells, such as young leaves, flowers, or shoots, to ensure a high abundance of expressed genes, though many samples also came from roots and other tissues.
+RNA extraction was performed using well-established protocols or commercial kits, facilitating the comprehensive analysis of transcribed RNA across this diverse set of species.
+This monumental effort not only sheds light on plant genetic diversity but also provides a rich data resource for ongoing and future research in plant science and evolutionary biology.
+We thank the reviewer’s constructive suggestion.
+In the camera-ready version, we will refine the broader limitations on the biological and ethical considerations, including the obstacles in in-vivo verifications, and implications in safety screening, etc.
+Further, we will add a section to discuss the broad potential and prospective applications (e.g., mRNA vaccine design) of our model and elaborate on the out-of-scope concepts and scenarios (e.g., tertiary structure prediction and generalization to other omics data).
+Appendix J Limitations
+Our work has certain limitations, which we are actively addressing in future iterations:
+Model and Data Scale: Although our current RNA FM outperforms prior arts in various scenarios, its scale remains small relative to the growing availability of large biological databases and the scaling laws described in Kaplan et al.
+(2020); Hoffmann et al.
+(2022); Muennighoff et al.
+(2023).
+Due to resource constraints, we were unable to pre-train substantially larger models or fully exploit large-scale databases like OneKP.
+Moving forward, we will focus on training more expansive foundation models to realize improved performance and generalization in both DNA and RNA contexts.
+Sequence Length Constraints: The modeling length of our current FMs, while adequate for many RNA and DNA tasks (given that RNA sequences are typically shorter than genomes), still may not be sufficient for some applications requiring very long-range modeling.
+Future directions will include enhancements to handle significantly longer sequences, enabling the model to tackle a wider variety of downstream tasks.
+Biological Verification and Ethical Considerations: A major challenge in the practical application of our models is the limited availability of cost-effective and ethically sound in vivo validation.
+The utility of predictions generated by our foundation models depends heavily on subsequent experimental studies, which are often expensive, time-consuming, and constrained by ethical and safety standards.
+Moreover, the use of these models in safety-critical scenarios, such as clinical diagnostics or therapeutic interventions, must be guided by rigorous safety screening, ethical oversight, and adherence to regulatory frameworks.
+Application Scope and Out-of-Scope Extensions: While we highlight prospective applications such as mRNA vaccine design and RNA-based therapeutics, our current approach does not directly address tasks like tertiary RNA structure prediction or generalization to other omics data types.
+Integrating these capabilities will require further methodological development, additional data modalities, and possibly more intricate architectures.
+Similarly, exploring other high-impact scenarios, such as broader gene-editing applications or regulatory RNA element identification, remains future work.
+Implications for Downstream Tasks: Although we have demonstrated encouraging results in several downstream applications, the complex functional landscapes of RNA biology mean that certain domain-specific tasks may still be challenging.
+To address these gaps, we plan to incorporate richer biological annotations, explore multi-modal inputs, and collaborate with domain experts to identify critical performance targets and safety checkpoints for biomedical use.
+Appendix K Ethics Statement
+This research utilised the open OneKP dataset, which is devoid of human-related privacy concerns.
+The pre-training sequences used are plant-based genomic data, which could have ecological implications.
+The following ethical guidelines must be adhered to when using OmniGenome:
+Ensure that the data is used responsibly, with proper attribution and fair compensation to the original sources.
+Prohibit the use of the model for unethical purposes, such as creating harmful bio-software or designing dangerous RNA structures.
+The models and findings should contribute to the conservation of plant species and their ecosystems, rather than posing a threat.
+Adhere to principles of transparency and open science, utilising publicly available datasets and providing comprehensive documentation of methodologies and findings.
+In conducting this research, we are committed to ethical scientific practices that respect biodiversity and contribute positively to genomic research.
+We advocate for ongoing dialogue regarding the ethical use of plant RNA sequences and support initiatives that ensure benefits from such research are shared equitably with all stakeholders.